find the area of a square if sides measure 24 cm
Found 2 solutions by rfer, chawki:
Answer by rfer(16322)

A=s^2
A=24^2
A=576 sq cm

Answer by chawki(7)

im not sure but if it says sides not side's, it means that the sum of all 4 sides are equal to 24, which means every side is 24/4 = 6cm
and thus the area would be 6*6 = 36